OpenStudy (anonymous):

subtract simplify by colecting like radical terms if posible. 2√(8)-2√(2)

OpenStudy (callisto):

2√(8)-2√(2) = 2√(2x2x2)-2√(2) = 2x2√(2)-2√(2) = 4√(2) - 2√(2) = 2√(2)

OpenStudy (anonymous):

thank you thats what i came up with as well

OpenStudy (callisto):

welcome :)